# CVPD at QIAS 2025 Shared Task: An Efficient Encoder-Based Approach for Islamic Inheritance Reasoning

### v1

@@ 1 Introduction
Large Language Models (LLMs) such as GPT-4 OpenAI (2023), Gemini Team et al. (2023), and Deepseek-v3 Liu et al. (2024) have advanced natural language processing, and show strong reasoning capabilities on many topics. However, as they were mainly trained on general web data, they often struggle in specialised domains with high accuracy Bubeck et al. (2023). Islamic inheritance law (ʿIlm al-Mawārīth) is one such area, which is based on fixed rules from the Qur’an and Sunnah and requires a precise understanding of the law and accurate mathematical proportion calculations Esmaeili (2012); Phillips and Wilson (1995). The complexity arises from rules such as farā'iḍ (fixed shares), ʿawl (reduction of shares if more than one), and radd (increase of shares if less than one) El-Far (2011), where errors can cause serious legal and financial problems. General LLMs often fail at such tasks due to the multi-step reasoning and strict numerical precision, especially in Arabic contexts Arabi and Hassan (2023). Reinforcing this point, a recent comprehensive study by Bouchekif et al. specifically assessed LLMs on Islamic inheritance law, providing empirical evidence of their limitations in this domain. Therefore, the QIAS 2025 SubTask 1 becomes a valuable benchmark for the assessment Bouchekif et al. (2025). This paper presents a lightweight framework developed for Islamic inheritance reasoning to address these challenges. Our approach combines a pre-trained Arabic text encoder with an Attentive Relevance Scoring (ARS) module. Instead of generating step-by-step generative answers, the system measures how strongly each possible answer relates to the question. The ARS module then ranks the options and selects the correct legal and mathematical outcome. This design focuses on accuracy and efficiency, providing a more feasible solution than large LLMs requiring high computational resources. We compare our specialised model with several leading general-purpose LLMs, including Gemini and DeepSeek, using the official QIAS 2025 dataset. Our experiments show that large models are prone to certain types of errors, especially under specific inference conditions. Our targeted approach, while not perfect, presents an alternative with a different performance and error profile, prioritizing consistency and efficiency. The primary contributions of this work are threefold:
@@ 4 Conclusion
We presented a lightweight framework for automated Islamic inheritance reasoning (ʿIlm al-Mawārīth), combining a specialized Arabic text encoder with an Attentive Relevance Scoring (ARS) mechanism for multiple-choice questions. Compared to leading API-based LLMs such as Gemini and DeepSeek, our local model using MARBERT achieved lower accuracy (69.87% vs. up to 87.60%) but offers significant advantages in efficiency, deployability, and data privacy, making it suitable for on-device applications.
These results highlight the trade-off between peak performance and practical usability, and demonstrate that lightweight, domain-adapted models remain valuable for high-stakes, rule-based tasks. Future work will aim to close the accuracy gap by integrating symbolic reasoning for inheritance calculations and exploring hybrid approaches that combine lightweight and large models for optimal performance and efficiency.

### v2

@@ 1 Introduction
Large Language Models (LLMs) such as GPT-4 OpenAI (2023), Gemini Team et al. (2023), and Deepseek-v3 Liu et al. (2024) have advanced natural language processing, and show strong reasoning capabilities on many topics. However, as they were mainly trained on general web data, they often struggle in specialised domains with high accuracy Bubeck et al. (2023). Islamic inheritance law (ʿIlm al-Mawārīth) is one such area, which is based on fixed rules from the Qur’an and Sunnah and requires a precise understanding of the law and accurate mathematical proportion calculations Esmaeili (2012); Phillips and Wilson (1995). The complexity arises from rules such as farā'iḍ (fixed shares), ʿawl (reduction of shares if more than one), and radd (increase of shares if less than one) El-Far (2011), where errors can cause serious legal and financial problems. General LLMs often fail at such tasks due to the multi-step reasoning and strict numerical precision, especially in Arabic contexts Arabi and Hassan (2023). Reinforcing this point, a recent comprehensive study by Bouchekif et al. (2025b) specifically assessed LLMs on Islamic inheritance law, providing empirical evidence of their limitations in this domain. Therefore, the QIAS 2025 SubTask 1 becomes a valuable benchmark for the assessment Bouchekif et al. (2025a). This paper presents a lightweight framework developed for Islamic inheritance reasoning to address these challenges. Our approach combines a pre-trained Arabic text encoder with an Attentive Relevance Scoring (ARS) module. Instead of generating step-by-step generative answers, the system measures how strongly each possible answer relates to the question. The ARS module then ranks the options and selects the correct legal and mathematical outcome. This design focuses on accuracy and efficiency, providing a more feasible solution than large LLMs requiring high computational resources. We compare our specialised model with several leading general-purpose LLMs, including Gemini and DeepSeek, using the official QIAS 2025 dataset. Our experiments show that large models are prone to certain types of errors, especially under specific inference conditions. Our targeted approach, while not perfect, presents an alternative with a different performance and error profile, prioritizing consistency and efficiency. The primary contributions of this work are threefold:
@@ 4 Conclusion
We presented a lightweight framework for automated Islamic inheritance reasoning (ʿIlm al-Mawārīth), combining a specialized Arabic text encoder with an Attentive Relevance Scoring (ARS) mechanism for multiple-choice questions. Our local model, using MARBERT, achieved a test accuracy of 69.87%, which, while promising, is notably lower than the 87.60% reached by leading API-based LLMs like Gemini. This performance difference stems from our model’s core design, which forgoes explicit, step-by-step symbolic reasoning in favor of efficient semantic matching. Despite this accuracy trade-off, our approach offers significant advantages in computational efficiency, on-device deployability, and data privacy, making it a viable solution for resource-constrained or offline applications.
These results highlight a critical trade-off between peak performance and practical usability. The current accuracy level positions our system as a valuable assistive tool for legal experts rather than a fully autonomous decision-maker, underscoring the necessity of human oversight in such high-stakes, rule-based domains. This demonstrates that lightweight, domain-adapted models remain highly relevant for specific use cases. Future work will directly aim to close the accuracy gap by integrating symbolic reasoning capabilities to handle the precise calculations inherent in inheritance law. We will also explore hybrid approaches that combine the efficiency of our lightweight model with the reasoning power of large models to achieve an optimal balance of performance and practicality.
Acknowledgement
The authors thank Mr. Arturo Argentieri from CNR-ISASI Italy for his technical contribution to the multi-GPU computing facilities. This research was partially funded by the Italian Ministry of University and Research (MUR) with the project "Future Artificial Intelligence Research —FAIR" Grant number PE0000013 CUP B53C22003630006.